OpenStudy (anonymous):

Log differentiation: Compute the following derivatives. Use log differentiation where appropriate. d/dx[(lnx)^(x^2)]

OpenStudy (anonymous):

original function: \[d/dx(lnx)^{x^2}\] Add ln on both sides \[y=(lnx)^{x^2}=>\ln(y)=\ln((lnx)^{x^2})\] use the log prop ln*x^r)=rln(x) on the right side: \[=>\ln(y)=\ln(x^2\ln(x))\] Take derivative using implicit diff. and remove the y in the denom \[=>y'/y=y(x) \times (\frac{ 1 }{ x^2\ln(x) })[(2x)(\ln(x))+(x^2)(\frac{1}{x})]\] Plug in y(x): \[y' = (lnx)^{x^2}[(\frac{1}{x^2\ln(x)})[(2x)(\ln(x)+(x^2)(\frac{1}{x})]\]

OpenStudy (anonymous):

The step with the log property is a bit off: \[\ln f(x)^{(g(x)}=g(x)\ln f(x)\] so \[\ln\left(\ln x\right)^{x^2}=x^2\ln(\ln x)\]

OpenStudy (anonymous):

@SithsAndGiggles I see so making those adjustments my new answer is: \[\ln(y)=(x^2)(\ln(\ln(x)))\] \[=>y'/y=y(x) \times(2x)(\ln(\ln(x))+\frac{1}{\ln(x)}* (\frac{1}{x})*(x^2)\] \[=>y'=[\ln(x)^{x^2}][(2x)(\ln(\ln(x))+(1/\ln(x)\times 1/x)\times x^2]\]

OpenStudy (anonymous):

Look good to me, just simplify 1/x*x^2 at the last term. :)

OpenStudy (anonymous):

ok, thanks.